OpenStudy (anonymous):

What is the equation of the line pretriceing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (anonymous):

what does pretrice mean?

OpenStudy (anonymous):

if it's perpendicular then the slope is 1/3

OpenStudy (anonymous):

by taking the negative reciprocal

OpenStudy (anonymous):

y - 1 = (1/3)(x - (-6) "using point slope" y - y1 = m(x - x1)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y - 1 = (1/3)(x - (-6) y=(1/3)x+2+1 y=(1/3)x+3